if a,b and c are three different complex numbers such that a/1-a=b/1-b=c/1-c=k then k=?

Grade:12th pass

1 Answers

Aditya Gupta
3 years ago
We see that x/(1-x)= k has roots a, b and c.
Or x(1+k) - k= 0 has more than one root which is possible only if 1+k= -k= 0
Or k= 0= -1, which is impossible.
So no possible value of k exists.
I am pretty sure that there is some mistake in the question, so you may repost a new question preferably with a picture.
KINDLY APPROVE :))